Transfer Pricing, Idle Capacity Mouton & Perrier, Inc., has a number of divisions that produce liquors, bottled water, and glassware. The Glassware Division manufactures a variety of bottles that can be sold externally (to soft-drink and juice bottlers) or internally to Mouton & Perriers Bottled Wat Division. Sales and cost data on a case of 24 basic 12-ounce bottles are as follows Unit selling price Unit variable cost Unit product fixed cost* Practical capacity in cases $350,000/500,000 During the coming year, the Glassware Division expects to sell 390,000 cases of this bottle. The Bottled Water Division currently plans to buy 100,000 cases on the outside market for $2.95 each. Ellyn Burridge, manager of the Glassware Division, approached Justin Thomas, manager of the Bottled Water Division, and offered to sell the 100,000 cases for $2.89 each. Ellyn explained to Justin that she can avoid selling costs of $0.12 per case by selling internally and that she would split the savings by offering a $0.06 discount on the usual price $2.95 $1.25 $0.70 500,000 Required 1. What is the minimum transfer price that the Glassware Division would be willing to accept? Round to the nearest cent. per unit What is the maximum transfer price that the Bottled Water Division would be willing to pay? Round to the nearest cent. per unit Should an internal transfer take place? Yes What would be the benefit (or loss) to the firm as a whole if the internal transfer takes place? Benefit V $ 2. Suppose Justin knows that the Glassware Division has idle capacity. Do you think that he would agree to the transfer price of $2.89? No Suppose he counters with an offer to pay $2.40. If you were Ellyn, would you be interested in this price? Yes 3. Suppose that Mouton & Perriers policy is that all internal transfers take place at full manufacturing cost. What would the transfer price be? Round to the nearest cent. per unit
Answer #1

Minimum transfer would be accept by Glassware Division = Variable Cost = $ 1.25

Maximum transfer would be willing to pay by Bottle water Division =Outside Market price = $ 2.95

Yes Internal trasfer should be take place from Glassware to Bottle water division as there is a idle capacity of 110000 units.

If Company purchase from outside then cost would be = no of units * Market price = 100000* 2.95 = $ 2950000

If Company procures from internal division the cost would be = no of units * variable cost = 100000* 1.25 = $125000

Net Benefit to the firm = 295000- 125000 = $170000

Note : Fixed cost wouldn't be consider if firm make additional 100000 units because Fixed cost would be same if firm make either 390000 units or upto 500000 units

2. No, Justin would not be agree because at the idle capacity cost of production to the glassware division will only be $ 1.25 i.e. variable cost. as fixed cost will remain same upto 500000 units

Yes, we would be interested because cost of production to the glassware division will only be $ 1.25 i.e. variable cost.

3. Manufacturing cost = variable cost + fixed cost per unit

= 1.25 + 0.70 = $1.95